OpenStudy (anonymous):

use the square root property please show the steps on getting the answer. (x−3)2−1=5

OpenStudy (mathmale):

(x−3)2−1=5 NEEDS to be re-written as (x−3)^2−1=5. Not to be picky, y ou understand, but for the sake of accuracy. Expand that (x-3)^2, please.

OpenStudy (mathstudent55):

@mathmale Isn't the question as king for this method? If \(X^2 = k\), then X\( = \pm \sqrt k\) where X = an expression in x.

OpenStudy (mathmale):

Indeed it is. But I wanted to get that 2nd power of the binomail correctly expanded.

OpenStudy (mathstudent55):

Don't expand the binomial. Instead, isolate it and take square roots of both sides.

OpenStudy (mathmale):

Have to agree with you. Simplify the given equation. How would you do that?

OpenStudy (mathstudent55):

Is that what is meant by "use the square root property"?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

How would you simplify (x−3)^2−1=5 ? forget about that (x-3)^2 for the moment.